An AC current with 50hz following sine law has arms value of 50 A. Write down the equation for instantaneous value calculate the instantaneous current 0.0025 s passing through positive maximum value.
 

Grade:7

2 Answers

Vikas TU
4 years ago
Dear student 
The equation od AC current will be 
I = I0 Sin(wt+Q)
Here I0 = 50 A 
f = 50 Hz 
w = 100*pie
I = 50 Sin(100*pie t + Q ) 
This will be the equation of current.
Good Luck 
Cheers 
Arun
4 years ago
f = 50 Hz I = I base 0 Sin Wt Peak value I = I base 0/√2 I base 0/√2 = I base 0 Sin Wt ⇒ I base 0/√2 Sin Wt = Sin π/4 ⇒ π/4 = Wt. or, t = π/400 = π/4 * 2πf = 1/8f = 1/8 * 50 = 0.0025 s = 2.5 ms